OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point(4,-5)

OpenStudy (anonymous):

slope intercept form : y = mx + b slope(m) = -3 (4,-5)...x = 4 and y = -5 now we use y = mx + b and sub in our stuff to find b, the y intercept. -5 = -3(4) + b -5 = -12 + b -5 + 12 = b 7 = b so your equation is : y = -3x + 7 any questions ?